9x16=wx12
my worksheet says its wrong but this it what i got...
9x16=144
144x12=1728
w=1,728

Answer by Alan3354(69443)

9x16=wx12
my worksheet says its wrong but this it what i got...
9x16=144
144x12=1728
-------------
Don't use x for multiply, use the asterisk.
9*16=w*12
144 = w*12
w = 144/12
w = 12